OpenStudy (anonymous):

Anna's weekly salary is $500 but she takes home only 395 dollars what percent of her weekly salary does she take home

OpenStudy (anonymous):

\[\frac{ 395 }{ 500}=\frac{ ? }{ 100}\] if you divide 500 to get 5 then you must divide 395 by 5 to get your answer 395\[\div5=79\] \[\frac{ 79 }{ 100}= 79%\]

OpenStudy (anonymous):

79 percent